# 1]Univ. Artois, UR 4025, Laboratoire Systèmes Electrotechniques et Environnement (LSEE), F-62400 Béthune, France 2]Univ. Artois, UR 3926, Laboratoire de Génie Informatique et d’Automatique de l’Artois (LGI2A), F-62400 Béthune, France Time series classification with random convolution kernels based transforms: pooling operators and input representations matter

### v1

@@ Abstract
This article presents a new approach based on MiniRocket, called SelF-Rocket, for fast time series classification (TSC). Unlike existing approaches based on random convolution kernels, it dynamically selects the best couple of input representations and pooling operator during the training process. SelF-Rocket achieves state-of-the-art accuracy on the University of California Riverside (UCR) TSC benchmark datasets.
@@
Random convolution kernels based transforms, ROCKET [1] and its successors, MiniRocket [2], MultiRocket [3] and others, have revolutionized the field of time series classification, offering an unrivaled compromise between speed and accuracy [4].
@@
In this article, a new approach based on MiniRocket, called SelF-Rocket, is developed. Unlike existing ones based on random convolution kernels, it dynamically selects the best data input representations and pooling operator combination during the training process.
This article is organized as follows. The fundamentals of time series classification and the main random convolution kernels based methods are recalled in Section 2. A statistical analysis, conducted on the University of California Riverside (UCR) archive [5], is then presented in Section 3 to motivate this work. It shows that while PPV is a pooling operator of choice, in most cases it is not the best one. The method SelF-Rocket is then introduced in Section 4 and its performances on the UCR archive are analyzed in Section 5. Finally, Section 6 offers a conclusion with some perspectives.
2 Related works
2.1 Time series classification
Time series classification (TSC) is a type of supervised learning problem where the objective is to assign a class label to a given time series instance. A Time series data consist of sequences of observations collected at constant intervals over time. Formally, a time series can be represented as:
@@ 2.1 Time series classification
Many methods of various kind have been developed to address TSC problems. Traditional methods include distance-based techniques like Dynamic Time Warping (DTW) [6], which measure the similarity between time series. More recently, Deep Learning approaches such as Inceptiontime [7] have been proposed. Hybrid approaches like HIVE-COTE v2.0, combining multiple classifiers of different kind, are generally the most accurate, at a high computing cost [8]. For a recent general survey of TSC methods, the interested reader may refer to [4].
In the next section, we will focus on a specific type of TSC methods using random convolution kernels to extract features from time series data.
@@ 2.2.1 Main methods
For each time series, features are extracted, where is the number of kernels.
@@
Weights are randomly selected and then normalized, .
@@
(6) where indicates the indices of positive values and
@@ 2.2.2 Kernel pruning
The main drawback of methods based on random convolutions is the large number of kernels, which may prevent them from being integrated into memory-limited devices. Moreover, not all kernels contribute positively to model performance, resulting in unnecessary computational complexity [10].
Various approaches have therefore been proposed for pruning the kernels, keeping only those most important to the problem.
@@
Pre-training the original model at full capacity,
@@
Initializing a population of binary candidate state vectors where each vector element represents the active/inactive status of a kernel,
@@
Using a population-based optimization algorithm to evolve this population, aiming to minimize the number of active kernels while maximizing classification accuracy,
@@
Combining the total number of active kernels with the classification accuracy to determine the best state vector, which is then used to select the optimal kernels.
@@
POCKET [11] improves S-ROCKET by dividing the pruning problem in 2 stages:
@@
Applying dynamically varying penalties for group-level regularization to prune redundant kernels.
@@
Uses element-level regularization to refit a linear classifier with the remaining features.
Experimental results demonstrate that POCKET performs 11 times faster than S-ROCKET while being as accurate.
Detach-ROCKET [12] prunes kernels using Sequential Feature Detachment (SFD), a method to identify and remove non-essential features. SFD involves iterative detachment steps, where a fixed percentage of the current active features is removed at each step. Depending on the number of classes, one or several RIDGE classifiers are trained on the set of active features, and RIDGE regressions are made to learn coefficients associated with each feature allowing one to eliminate the least relevant features.
@@ 2.2.3 ROCKET ensembles
FT-FVC [13] adds feature diversity to ROCKET using three transformations to raw data: Hilbert and first-order and second-order difference. These transformations generate feature vectors that are concatenated with the raw feature vector to form a three-view representation. These representations are merged through hard voting.
@@ 3 On the importance of choosing the right input representations and pooling operator
To understand how input representations and pooling operators impact the performances of random convolution kernels based transforms, we compared, on the 112 datasets of the UCR archive [5], original MiniRocket with modified versions using GMP, MPV, MIPV and LSPV instead of PPV as pooling operator and the first order difference as additional input representation, leading to possible transforms (detailed in Section 4.1). In this article, transforms using multiple pooling operators were not considered since preliminary results showed they under-performed those using only one pooling operator in almost all the cases. Table 1 shows the mean classification accuracy across resamples of the possible transforms for some datasets included in UCR archive. For each dataset, the number in bold indicates the greatest overall accuracy achieved. Note that the PPV column is also the performance of the original MiniRocket classifier. Figure 1 shows the number of datasets in the UCR archive for which the modified MiniRocket version performs the best. If PPV_MIX is the best transform on average, it is outperformed by others in most cases ().
Figure 2 shows the critical difference diagram [14] of the 15 possible transforms for the different datasets, the distribution of classification accuracy is provided as box plots in Figure 13.
@@ 4 SelF-Rocket
Following the experimental results obtained in the previous section, we investigate a new variant of ROCKET, called SelF-Rocket (Selected Features Rocket), that aims at dynamically selecting the best couple of input representations and pooling operator during the training process. The latter involves three main stages:
SelF-Rocket works in a similar way to MiniRocket or MultiRocket. The only major difference between them is the addition of a feature selection phase. As seen previously in Section 3, the optimal set of features varies from one dataset to another, so using a single feature type (PPV) like MiniRocket or a fixed set of feature types ({PPV, LSPV, MIPV, MPV}) like MultiRocket can produce good results on average, but not optimal for each case. To address this problem, we implement a wrapper-based feature selection module with the objective to select the most appropriate set for each classification problem.
@@ 4.1 Feature Generation
Let IR be a set of Input Representations, PO a set of Pooling Operators, MK the set of all kernels (containing every kernel/dilation combinations) generated by MiniRocket and the bias associated to a kernel (cf Section 2.2.1).
@@
Let be a feature generation function defined as:
@@ 4.2 Feature Selection
The Feature Selection Module relies on the Stratified -Fold method to create novel train and validation sets derived from the original train set. These sets are then used to train mini-classifiers for each IR-PO combination. Subsequently, the optimal combination on average is selected as the final set of features for the linear classifier. The Feature Selection Module is summarized in Algorithm 1.
Training set class y_train
The Number of folds k
The Number of features per mini-classifier f
@@ 5 Experiments
In this section, the performances of SelF-Rocket are evaluated.
We show that SelF-Rocket is as accurate as Hydra-MultiRocket and HIVE-COTE v2.0, despite having only one type of feature, and also has a relatively shorter classifier training time. We also investigate the
influence of the main parameters on the model performance.
@@ 5.1 Experimental settings
SelF-Rocket has been evaluated on a subset of the UCR Time Series Archive [5], which is widely used in this field. To guarantee the reproducibility of our experiments and a fair comparison with other TSC algorithms, we employed the identical 112 datasets as those used by [4] and the precise same 30 train/test resamples for each of those datasets. The 112 datasets are those from the 128 original ones that did not contain missing values or variable time series length.
Note that we did not compute the results for others methods shown in Figures 4, 5 and 6, but used the results111https://github.com/time-series-machine-learning/tsml-eval/tree/main/results/classification/Univariate available in [4].
The author’s implementation of MiniRocket222https://github.com/angus924/minirocket/tree/main was used as baseline model. Our algorithm is implemented in standard python 3.11.4.
In order to compare the classification performance with other TSC algorithms, we use the critical difference diagrams that display mean ranks of each method across all datasets with the horizontal cliques indicating that there is no statistically significant difference between those methods, the Multiple Comparison Matrix (MCM) [15] with heat-map color representing mean differences in score, and the pairwise scatter plots of test accuracy summarizing the number of win/draw/loss between 2 classifiers across all datasets.
All experiments were conducted on a Dell Precision 5480 on Debian 12 with an Intel Core i9-13900H vPro 2.6 GHz and 32GB of RAM
@@ 5.2 Comparison with other TSC methods
We compare the performances of SelF-Rocket with 11 TSC methods: 1NN-DTW, Elastic Ensemble (EE) [16], Shapelet transform classifier (STC) [17], HIVE-COTE v2.0 [8], H-Inceptiontime [18], MultiRocket [3], Hydra, Hydra+MultiRocket [9], MiniRocket [2],
@@
Figure 4 shows the average mean accuracy rank of these classifiers. In terms of performance, SelF-Rocket ranks second only to HIVE-COTE v2.0, and outperforms others.
Figure 5 shows the Multiple Comparison Matrix [20], i.e. the full pairwise comparison between the best performing methods tested. The mean difference in accuracy between SelF-Rocket and its baseline MiniRocket is approximately 1.23 %
Figure 6 displays the mean accuracy, the pairwise win/draw/loss, and value for statistical tests between the tested methods over 30 resamples for the 112 selected UCR datasets. Each point represents a dataset, the more a point is distant to the diagonal line, the more one of the two methods performs better than the other. In some cases, SelF-Rocket outperforms Hydra-MultiRocket, but in general, both methods offer comparable performance. HIVE-COTE v2.0 also has the advantage over SelF-Rocket in certain datasets.
There is no significant statistical difference by Wilcoxon signed rank test between SelF-Rocket, Hydra-MultiRocket and HIVE-COTE v2.0
SelF-Rocket (9,996 or 19,992 features) is faster in the training and prediction phase than MultiRocket and Hydra-MultiRocket due to a lower number of features (resp 50,000 and 50,000 + number of features of Hydra).

### v4

# 1]Univ. Artois, UR 4025, Laboratoire Systèmes Electrotechniques et Environnement (LSEE), F-62400 Béthune, France 2]Univ. Artois, UR 3926, Laboratoire de Génie Informatique et d’Automatique de l’Artois (LGI2A), F-62400 Béthune, France Time series classification with random convolution kernels: pooling operators and input representations matter

@@ Abstract
This article presents a new approach based on MiniRocket, called SelF-Rocket (Selected Features Rocket), for fast time series classification (TSC). Unlike existing approaches based on random convolution kernels, it dynamically attempts to select the best couple of input representations and pooling operator during the training process. SelF-Rocket achieves state-of-the-art accuracy on the University of California Riverside (UCR) TSC benchmark datasets.
@@
Random convolution kernels based transforms, ROCKET [1] and its successors, MiniRocket [2], MultiRocket [3] and others, have significantly contributed to the advancement of time series classification, offering an unrivaled compromise between speed and accuracy [4].
@@
In this article, a new approach based on MiniRocket, called SelF-Rocket (Selected Features Rocket), is developed. Unlike existing ones based on random convolution kernels, it dynamically attempts to select the best couple (or combination) of data input representations and pooling operator during the training process. The complete source code of our implementation of SelF-Rocket is available on GitHub111https://github.com/ANR-MYEL/SelF-Rocket/.
This article is organized as follows. The fundamentals of time series classification and the main random convolution kernels based methods are first recalled in Section 2. A statistical analysis, conducted on the University of California Riverside (UCR) archive [5], is then presented in Section 3 to motivate this work. It shows that while PPV is a pooling operator of choice, in most cases it is not the best one. The method SelF-Rocket is then introduced in Section 4 and its performances on the UCR archive are analyzed in Section 5. Finally, Section 6 gives a conclusion with some perspectives.
2 Time series classification
2.1 Overview
Time Series Classification (TSC) is a type of supervised learning problem where the objective is to assign a class label to a given time series instance. The data in a time series is made up of sequences of observations collected at constant intervals over time. Formally, a time series can be represented as:
@@ 2.1 Overview
Many methods of various kinds have been developed to address TSC problems. Traditional methods include distance-based techniques like Dynamic Time Warping (DTW) [6], which measure the similarity between time series. More recently, Deep Learning approaches such as Inceptiontime [7] have been proposed. Hybrid approaches like HIVE-COTE v2.0, combining multiple classifiers of different kinds, are generally the most accurate, at a high computing cost [8]. For a recent general survey of TSC methods, the interested reader may refer to [4].
The next section focuses on a particular type of CST method that involves the use of random convolution kernels for feature extraction from time series data.
@@ 2.2.1 Main methods
For each kernel, two features are extracted using PPV and GMP.
@@
Weights are randomly selected and then normalized: .
@@
(6) where indicates the indices of positive values, and
@@ 2.2.2 Kernel pruning
The primary limitation of methods that rely on random convolutions is the extensive number of kernels they require, which may prevent them from being integrated into memory-limited devices. Moreover, not all kernels contribute positively to model performance, resulting in unnecessary computational complexity [10].
As a result, several methods have been suggested for pruning the kernels, retaining only those that are most relevant to the problem.
@@
The original model is initially pre-trained to its full capacity.
@@
A population of binary candidate state vectors is initialized, where each element in a vector indicates whether a kernel is active or inactive.
@@
A population-based optimization algorithm is employed to evolve this population, with the goal of reducing the number of active kernels while enhancing classification accuracy.
@@
The total number of active kernels and the classification accuracy are combined to identify the best state vector, which is then used to select the optimal set of kernels.
@@
POCKET [11] enhances S-ROCKET by addressing the pruning problem in two distinct stages:
@@
It applies dynamically varying penalties for group-level regularization to remove redundant kernels.
@@
It employs element-level regularization to refit a linear classifier using the remaining features.
Experimental results show that POCKET performs 11 times faster than S-ROCKET while being as accurate.
Detach-ROCKET [12] prunes kernels using Sequential Feature Detachment (SFD), a technique to identify and remove non-essential features. SFD involves iterative detachment steps, where a fixed percentage of the current active features is removed at each step. Depending on the number of classes, one or several Ridge classifiers are trained on the set of active features, and Ridge regressions are made to learn coefficients associated with each feature allowing one to eliminate the least relevant features.
@@ 2.2.3 ROCKET ensembles
FT-FVC [13] adds feature diversity to ROCKET using three transformations to raw data: Hilbert, first-order and second-order difference. These transformations generate feature vectors that are concatenated with the raw feature vector to form a three-view representation. Three classifiers are trained separately on each view, and a prediction is made by hard voting.
@@ 3 On the importance of choosing the right input representations and pooling operator
To assess the influence of input representations and pooling operators on the performance of transformations based on random convolution kernels, we conducted a comparison using 112 selected datasets of the UCR archive [5]. This comparison involved the original MiniRocket and modified versions, which utilized GMP, MPV, MIPV, and LSPV as pooling operators instead of PPV. Additionally, we incorporated the first-order difference (DIFF) as an extra input representation. Classifiers are denoted using a notation PO_IR depending on the pooling operator PO used () and the input representation IR used (MIX denoting the concatenation of the first order difference DIFF and the base representation). We then compared possible transformations, which are formally defined in Section 4.1. These transformations correspond to combinations of these 5 potential pooling operators with these 3 possible input representations: the base representation, Diff, and MIX. It is worth mentioning that we also conducted experiments using multiple pooling operators, rather than selecting just one as described here. However, our preliminary results suggested that using multiple operators did not yield performances as effective as those discussed in this section.
Table 1 illustrates the mean classification accuracy across resamples of some datasets included in UCR archive for the possible transforms. For each dataset, the number in bold indicates the highest overall accuracy achieved. Note that the PPV column is also the performance of the original MiniRocket classifier.
Figure 1 shows the number of datasets in the UCR archive for which the modified MiniRocket version performs the best. If PPV_MIX is the best transform on average, it is outperformed by others in most cases ().
Figure 2 displays the critical difference diagram [14] of the 15 possible transforms for the different datasets. In Appendix A, the distribution of classification accuracy is also provided as box plots in Figure 18.
@@
Based on these experiments, it is evident that the combination of input representations and pooling operators yielding the best performance varies depending on the dataset. Therefore, it would be beneficial to have the capability to select the appropriate combination based on a training dataset. This is the goal of the SelF-Rocket method, which will be presented in the following section.
@@ 4 SelF-Rocket
Following the experimental results obtained in the previous section, we investigate a new variant of ROCKET, called SelF-Rocket (Selected Features Rocket), that aims to dynamically select the best couple of input representations and pooling operator during the training process. The latter, illustrated in Figure 3, involves three main stages:
SelF-Rocket operates similarly to MiniRocket or MultiRocket, the main difference being the inclusion of a feature set selection step. As previously observed in Section 3, the optimal set of features can differ from one dataset to another. Therefore, while using a single feature type like PPV, as in MiniRocket, or a fixed set of feature types such as {PPV, LSPV, MIPV, MPV}, as in MultiRocket, can yield generally good results, it may not be optimal for every specific case. To tackle this problem, we have incorporated a wrapper-based Feature Selection Module with the objective to select the most appropriate set of features for each classification problem.
@@ 4.1 Feature Generation
Let IR be a set of Input Representations, PO a set of Pooling Operators, MK the set of all kernels (containing every kernel/dilation combination) generated by MiniRocket and the bias associated to a kernel (cf Section 2.2.1).
@@
Let be a feature set generation function defined as:
@@ 4.2 Feature Selection
Summarized in Figure 4 and Algorithm 1, the Feature Selection Module relies on stratified train/test split methods to generate new train and validation sets derived from the original train set. Depending on the size of that dataset, we use either a stratified -fold repeated times or a stratified shuffle split with splits. A stratified shuffle split is used if the dataset contains at least a sufficiently large number of data (denoted by in Algorithm 1), it should be noted that this value has a direct impact on the time available for finding the best IR-PO combination by limiting the maximum size of each train and validation set, and thus the time available for this task. This strategy ensures that, as the size of the dataset increases, the learning and validation sets will vary more significantly from one voter to another. It therefore allows more reliable assessments of the performance of the different combinations. These sets are then used to train mini-classifiers for each IR-PO combination.
Subsequently, the selection of the optimal IR-PO combination across all the original train set splits is made using a highest median voting system, which is validated by the Algorithm 2 prior to its use as the final set of features for the linear classifier. Using the highest median to select an IR-PO as an alternative to the highest mean is preferable as it is more robust to extreme values (e.g. lucky runs) that may occur. The idea of the Algorithm 2 is to check if the selected IR-PO combination is sufficiently supported by the voters, otherwise a default combination is chosen. Its aim is to avoid choosing an IR-PO combination that has poor generalisability, especially with small datasets. In our implementation, Algorithm 2 checks if the selected IR-PO combination is part of the top values for each voter (e.g. in the top for each voter) with a certain degree of flexibility (e.g. at , implemented as threshold in Algorithm 2), otherwise the selected IR-PO combination is replaced by a default one (PPV_MIX in our case).
Training set class
The number of features per mini-classifier
The number of runs
The maximum dataset size
Index of the chosen IR-PO combination
The number of voters
The top considered
The value of the threshold
@@ 5 Experiments
In this section, the performances of SelF-Rocket and Hydra + SelF-Rocket (concatenation of the features of Hydra and SelF-Rocket) are evaluated. We show that SelF-Rocket is as accurate as MultiRocket, despite the use of only one pooling operator at a time, and also has a relatively shorter classifier prediction time. We also investigate the influence of the main parameters of SelF-Rocket on its performance.
@@ 5.1 Experimental settings
To guarantee the reproducibility of our experiments and a fair comparison with other TSC algorithms, the proposed method is evaluated on the same same identical 112 datasets of the UCR Time Series Archive [5] and the precise same 30 train/test resamples for each of those datasets [1, 9, 15, 3, 16, 4]. The 112 datasets are those from the 128 original ones that did not contain missing values or variable time series length.
The evaluation of a TSC method on UCR may be limited [5, 17] (for example, the data are pre-processed and of the same length), nevertheless, as described above, it currently remains the standard for benchmarking TSC methods.
Note that we did not compute the results for other methods shown in Figures 5, 6 and 7, but used the results222https://github.com/time-series-machine-learning/tsml-eval/tree/main/results/classification/Univariate available in [4].
The original implementations of MiniRocket333https://github.com/angus924/minirocket and MultiRocket444https://github.com/ChangWeiTan/MultiRocket were used as baseline. Our algorithm is implemented in standard Python 3.11.4.
In order to compare the classification performance with other TSC algorithms, we use the critical difference diagrams that display mean ranks of each method across all datasets with the horizontal cliques indicating that there is no statistically significant difference between those methods, the Multiple Comparison Matrix (MCM) [18] with heatmap color representing mean differences in score, and the pairwise scatter plots of test accuracy summarizing the number of win/draw/loss between 2 classifiers across all datasets.
All experiments were conducted on a cluster using Ubuntu 24.04.1 LTS with an Intel(R) Xeon(R) Gold 6434 and 250GB of RAM.
A sensitivity analysis of the main parameters of Self-Rocket is given in Section 5.3. We have chosen as default parameters for SelF-Rocket, a number of folds of , a number of features per mini-classifier of , a number of runs of , and a validation vote with a top and a threshold of . The others parameters of SelF-Rocket, i.e. number of kernels, padding, dilation, bias, length of kernel, values inside the kernel, remain the same as the default parameters of MiniRocket.
@@ 5.2 Comparison with other TSC methods
We compare the performances of SelF-Rocket and Hydra + SelF-Rocket with 11 TSC methods: 1NN-DTW, Elastic Ensemble (EE) [19], Shapelet transform classifier (STC) [20], HIVE-COTE v2.0 [8], H-Inceptiontime [21], MultiRocket [3], Hydra, Hydra+MultiRocket [9], MiniRocket [2],
@@
| Mean-Accuracy | HC2 0.8914 | H + SelF-R 0.8871 | H + MR 0.8840 | SelF-R 0.8832 | MultiR 0.8814 | H-IT 0.8761 | MiniR 0.8744 | QUANT 0.8671 |
| HC2 0.8914 | Mean-Difference rc / r=c / rc Wilcoxon p-value | 0.0043 60 / 3 / 49 0.2374 | 0.0074 59 / 7 / 46 0.1046 | 0.0082 73 / 3 / 36 0.0001 | 0.0100 67 / 6 / 39 0.0051 | 0.0153 73 / 4 / 35 0.0007 | 0.0170 84 / 5 / 23 1e-04 | 0.0243 81 / 5 / 26 1e-04 |
| H + SelF-R 0.8871 | -0.0043 49 / 3 / 60 0.2374 | - | 0.0031 56 / 4 / 52 0.4733 | 0.0039 75 / 4 / 33 1e-04 | 0.0056 64 / 4 / 44 0.0212 | 0.0110 68 / 3 / 41 0.0061 | 0.0127 87 / 4 / 21 1e-04 | 0.0200 77 / 4 / 31 1e-04 |
| H + MR 0.8840 | -0.0074 46 / 7 / 59 0.1046 | -0.0031 52 / 4 / 56 0.4733 | - | 0.0008 67 / 5 / 40 0.0015 | 0.0026 68 / 9 / 35 0.0003 | 0.0079 66 / 4 / 42 0.0644 | 0.0096 80 / 5 / 27 1e-04 | 0.0169 75 / 5 / 32 1e-04 |
| SelF-R 0.8832 | -0.0082 36 / 3 / 73 0.0001 | -0.0039 33 / 4 / 75 1e-04 | -0.0008 40 / 5 / 67 0.0015 | - | 0.0017 47 / 5 / 60 0.1561 | 0.0071 62 / 5 / 45 0.2066 | 0.0088 78 / 7 / 27 1e-04 | 0.0161 74 / 5 / 33 0.0001 |
| MultiR 0.8814 | -0.0100 39 / 6 / 67 0.0051 | -0.0056 44 / 4 / 64 0.0212 | -0.0026 35 / 9 / 68 0.0003 | -0.0017 60 / 5 / 47 0.1561 | - | 0.0054 63 / 4 / 45 0.1944 | 0.0070 80 / 5 / 27 1e-04 | 0.0143 75 / 5 / 32 1e-04 |
| H-IT 0.8761 | -0.0153 35 / 4 / 73 0.0007 | -0.0110 41 / 3 / 68 0.0061 | -0.0079 42 / 4 / 66 0.0644 | -0.0071 45 / 5 / 62 0.2066 | -0.0054 45 / 4 / 63 0.1944 | - | 0.0017 58 / 4 / 50 0.5043 | 0.0090 61 / 5 / 46 0.0502 |
| MiniR 0.8744 | -0.0170 23 / 5 / 84 1e-04 | -0.0127 21 / 4 / 87 1e-04 | -0.0096 27 / 5 / 80 1e-04 | -0.0088 27 / 7 / 78 1e-04 | -0.0070 27 / 5 / 80 1e-04 | -0.0017 50 / 4 / 58 0.5043 | - | 0.0073 60 / 5 / 47 0.0953 |
| QUANT 0.8671 | -0.0243 26 / 5 / 81 1e-04 | -0.0200 31 / 4 / 77 1e-04 | -0.0169 32 / 5 / 75 1e-04 | -0.0161 33 / 5 / 74 0.0001 | -0.0143 32 / 5 / 75 1e-04 | -0.0090 46 / 5 / 61 0.0502 | -0.0073 47 / 5 / 60 0.0953 | If in bold, then p-value 0.05 |
The average accuracy rank of these classifiers is displayed in Figure 5. In terms of performance, Hydra + SelF-Rocket ranks second only to HIVE-COTE v2.0, and outperforms others. SelF-Rocket ranks just after MultiRocket.
Figure 6 depicts the Multiple Comparison Matrix [22], i.e. the full pairwise comparison between the best performing methods tested. The mean difference in accuracy between SelF-Rocket and its baseline MiniRocket is approximately 0.88 %. We can also see that SelF-Rocket has a slightly higher accuracy than MultiRocket, despite having fewer wins by pairwise comparison.
Figure 7 displays the mean accuracy, the pairwise win/draw/loss, and value for statistical tests between the tested methods over 30 resamples for the 112 selected UCR datasets. Each point represents a dataset; the more a point is distant from the diagonal line, the more one of the two methods performs better than the other. In some cases, Hydra + SelF-Rocket outperforms Hydra + MultiRocket, but in general, both methods offer comparable performance. HIVE-COTE v2.0 also has the advantage over Hydra + SelF-Rocket in certain datasets.
There is no significant statistical difference by the Wilcoxon signed rank test between Hydra + SelF-Rocket, Hydra + MultiRocket, and HIVE-COTE v2.0.
SelF-Rocket (9,996 or 19,992 features) and Hydra + SelF-Rocket (9,996 or 19,992 features + number of features of Hydra) are faster in the classifier prediction phase than MultiRocket (50,000 features) and Hydra + MultiRocket (50,000 + number of features of Hydra) due to a lower number of features.
Figure 8 shows the IR-PO combinations selected from the 30 resamples in each of the 112 UCR data sets in three scenarios: before the vote validation step (cf Figure 4 and Algorithm 2), when only votes with high consensus are considered (at least 90% of voters must include the chosen combination in their top 5), and after the vote validation step (SelF-Rocket method). For the latter, we can see that several IR-PO combinations can be chosen, even if the PPV_MIX combination remains the most frequently chosen, mainly due to its use as the default combination when insufficient consensus is reached (cf Algorithm 2). Before the vote validation phase, several different IR-PO combinations could also be selected. However, experiments (cf 5.3.5) show suboptimal performance, hence the need for a vote validation phase to ensure the selection of a better IR-PO combination. Finally, the distribution of IR-PO combinations for votes resulting in high voter consensus (votes requiring readjustment are not considered here) is closest to the distribution of the best Oracle IR-PO combinations (cf Figure 1).